OpenStudy (anonymous):

Samuel currently has an account balance of $3,439.96. He opened the account five years ago with a deposit of $3,104.31. If the interest compounds quarterly, what is the interest rate on the account?

OpenStudy (anonymous):

3439.96=C(1+i/4)^20

OpenStudy (anonymous):

Try to solve for i and i'll tell you if you're correct.

OpenStudy (anonymous):

a) 4.6% b)8.3% c).5% d) 2.1% I did A=P(1+r/n)^nt I have up to 1.108= (1+x/4)^20 I don't know how to evaluate the exponent

OpenStudy (anonymous):

You do the 20th square of 1.108.

OpenStudy (anonymous):

Would that require logs? I don't really remember how to do that

OpenStudy (anonymous):

Which would be 1.005141.

OpenStudy (anonymous):

so 2.1% ?!